# Mohan Xu1, Kai Li1,2,, Guo Chen1,2 & Xiaolin Hu1,2,3,†

### v1

@@ 3.3.2 Full-frequency-frame attention module
$\boldsymbol{O}_{i}=\text{Softmax}\left(\frac{\boldsymbol{Q}_{i}\boldsymbol{K}_{i}^{\text{T}}}{\sqrt{E\times T}}\right)\boldsymbol{V}_{i},$ (7)
@@ 3.4 Band restoration module
After going through the separator, the sub-bands need to be converted back to their original width during mask estimation. Specifically, $\boldsymbol{J}\in\mathbb{R}^{N\times K\times T}$ denotes the output of the separator. For the $k$-th sub-band feature $\boldsymbol{J}_{k}\in\mathbb{R}^{N\times T}$ ($k\in[1,K]$), the PReLU activation function and 1D convolutions are used to transform the number of channels to twice the original dimension $2G_{k}$, corresponding to the real and the imaginary part. The complex feature is restored to generate a mask for each sub-band $\boldsymbol{M}_{k}\in\mathbb{C}^{G_{k}\times T}$ using the ReLU activation function. Then they are merged on the frequency dimension to get the mask for the whole band $\boldsymbol{M}\in\mathbb{C}^{F\times T}$. Similar to band-split, the 1D convolutions of different sub-bands do not share parameters.
@@ 4 EchoSet
SoundSpaces 2.0 is an audio rendering platform in 3D environments. Given the mesh of a 3D scenario, it can simulate the acoustic effects of any sound captured from microphones. We followed the steps below to generate mixed speech. (1) Choose the scenario. We selected rooms where daily conversations often occur (such as office, living room, bedroom, dining room, etc.) from Matterport3D, a large RGB-D dataset containing 90 diverse multi-floor and multi-room indoor scenes. (2) Define or sample the position. We defined a microphone at a suitable position, like next to a table or sofa, and sampled two sound sources in the same room. (3) Sample the direction. The angle between the microphone and the sound source must be obtuse, meaning that the speaker and listener face each other. (4) Sample the height. The microphone and sound sources were randomly generated at a vertical height of 1.5 m to 1.9 m from the floor, which is about a person’s height. (5) Generate the audio. With SoundSpaces 2.0, mixed audio files were generated based on bidirectional path tracking algorithm (Cao et al., 2016), which can simulate various effects in the sound propagation process, including reverberation, diffraction, and absorption. Materials of the room wall and the objects were annotated by Matterport3D and considered during the generation of the audio mixture.
Based on the SoundSpaces 2.0 platform and the Matterport 3D scene dataset, we can simulate reverberant audio from different speakers in LibriSpeech (Panayotov et al., 2015) to build a new dataset, EchoSet. In total, EchoSet includes 20,268 training utterances, 4,604 validation utterances, and 2,650 test utterances. Each utterance lasts for 6 seconds. We mixed the speech of the two speakers at a random overlap ratio and added some noises from WHAM! noise (Wichern et al., 2019). The two different speakers were mixed with signal-to-distortion ratio (SDR) sampled between -5 dB and 5 dB. The noises were mixed with SDR sampled between -10 dB and 10 dB.
@@ 5 Experimental Setup
Dataset. We report the performance of TIGER on EchoSet. For fair comparison with previous speech separation methods (Li et al., 2023; Wang et al., 2023; Hu et al., 2021), we also used two benchmark datasets LRS2-2Mix (Li et al., 2023) and Libri2Mix (Cosentino et al., 2020).
To validate the gap between EchoSet and real-world environments, we constructed real-world data by selecting 10 real-world environments and recording audio from 40 speakers from the LibriSpeech (Panayotov et al., 2015) test set. The two audio used for mixing were recorded in the same acoustic scene (e.g., the shape and material of the walls and objects in the room) and followed the same mixing method as LRS2-2Mix (Li et al., 2023). For more details of these datasets, please refer to Appendix A.
Training and evaluation. For training, adam optimizer (Kingma & Ba, 2014) was employed with an initial learning rate of 0.001, adjusted based on validation performance. Evaluation metrics included SDRi and SI-SDRi (Vincent et al., 2006), with higher values indicating better performance. We report parameters and MAC operations for complexity, which are calculated for one second of audio at 16 kHz. Inference speed was measured on NVIDIA RTX 4090 and Intel Xeon Gold 6326. Detailed training and evaluation configurations can be found in Appendix B and Appendix C. The code will be made publicly available after the paper is accepted.
@@ 6.1 EchoSet is more close to the real-world data
Figure 4: Test results of different models on the real-world data. Models were trained on LibriMix, LRS2-2Mix and EchoSet respectively.
@@ 6.2 Comparisons with state-of-the-art methods
Table 2: Performance comparison of TIGER and other separation models on Libri2Mix, LRS2-2Mix, and EchoSet. Models are trained and tested on corresponding datasets. Bold denotes the best performance, and underline indicates the second-best.
@@
Separation performance. TIGER obtained competitive separation performance on the three datasets compared with previous SOTA models (see Table 2). On Libri2Mix, which is relatively simple for lack of noise and reverberation, TIGER (large) was second only to the current SOTA model TF-GridNet, with a 6% drop in performance. On LRS2-2Mix, a more complicated dataset with reverberation recorded in different scenes, the drop in performance of TIGER (large) was only 2% compared with TF-GridNet. On EchoSet, the only dataset with the most realistic reverberation among the three, TIGER (large) achieved an SDRi of 14.22 dB, surpassing other existing methods. On this dataset, TIGER (small) also achieved the performance that was only slightly lower than TF-GridNet. From the above experimental results, we can see that the more complex the acoustic scenarios are, the better performance TIGER will produce. Similarly, based on the results in Figure 4, we observed that TIGER also outperforms existing models in real-world test scenarios. This demonstrates that TIGER is applicable to complex real-world acoustic scenarios including diverse noise and reverberation. To visualize the separation result, we present the spectrogram differences between the audio separated by TIGER and TF-GridNet (Appendix H), demonstrating that TIGER is capable of effectively reconstructing both low-frequency and high-frequency features.
Separation effciency. The parameters of TIGER were only 0.82 M, and the MACs were only 7.65 G/s and 15.27 G/s for the small and large versions respectively. Compared with TF-GridNet, the parameters of TIGER (large) dropped by 94.3%, and the MACs were reduced by 95.3%. For inference of one-second audio, TIGER (large) took about 1/3 of the CPU Time and 3/4 of the GPU Time compared with TF-GridNet, demonstrating a significant calculation compression effect. Besides, TIGER took up less memory during training and inference, making TIGER more suitable for deployment on devices with limited computational resources. To our knowledge, TIGER is the first neural network-based speech separation model with a parameter less than 1M.
Performance on cinematic sound separation. To demonstrate the generalization ability of TIGER on similar tasks, we migrate TIGER to cinematic sound separation, which aims to extract different audio elements from a film’s soundtrack. Compared with the previous SOTA model BSRNN, the parameters drop by 97.3% and MACs drop by 77.6%, with a 39.4% increase of SDR in performance. Details of the dataset, training configuration and result analysis can be found in Appendix D. We have provided audio and video demonstrations of the speech and cinematic sound separation. Please refer to the project page [^1] https://cslikai.cn/TIGER/.
@@ 6.3 Ablation study
We also verified the impact of alternating the frequency path and frame path on model performance (Appendix G). The results show that the time-frequency interleaved structure captures acoustic information of audio better than modeling only time or only frequency.
@@ Appendix A Dataset details
EchoSet. This dataset includes 20268 training utterances, 4604 validation and 2650 test utterances. The length of each audio is 6 seconds. The target speech was selected from LibriSpeech  (Panayotov et al., 2015), mixed with SDR ranging from -5 dB to 5 dB. The speech and noise which was sampled from WHAM! were mixed with SDR sampled between -10 dB and 10 dB. This dataset contains realistic reverberation.
LRS2-2Mix (Li et al., 2023). Each audio in this dataset lasts for 2 seconds, and the training set, validation set and test set are about 11.1, 2.8 and 1.7 hours, respectively. The utterances were selected from the LRS2 (Afouras et al., 2018) corpus, which consists of video clips acquired through BBC, and were mixed with SDR sampled between -5 dB and 5 dB. Since the audio files were recorded in real acoustic scenarios, LRS2-2Mix contains much noise and reverberation.
Libri2Mix (Cosentino et al., 2020). Each audio in this dataset lasts for 3 seconds. The target speech for each audio mixture was randomly chosen from LibriSpeech (Panayotov et al., 2015) (train-100) and combined with a uniformly sampled Loudness Units relative to Full Scale (Series, 2011) between -25 and -33 dB. This dataset contains no noise or reverberation.
Real-world data. We collected a smaller-scale real-world dataset to test the performance of models trained on different datasets. First, we randomly selected audio from various speakers in the LibriSpeech (Panayotov et al., 2015) test set, totaling 1.5 hours. All of this audio was clean, without reverberation or noise. The noise data were selected from the WHAM! noise dataset (Wichern et al., 2019). Next, we recorded the selected audio in 10 real-world environments, primarily consisting of rectangular conference rooms of various sizes and rooms with different shapes (such as fan-shaped and circular conference rooms). Finally, we mixed the audio from different speakers, which moved in different trajectories using SDRs sampled between -5 dB and 5 dB, while the noise was mixed using SDRs sampled between -10 dB and 10 dB.
@@ Appendix B Training Configuration
During training, We used a 3-second audio segment for EchoSet and Libri2Mix, and a 2-second for LRS2-2Mix. We used the maximization of SI-SDR as the training loss (Li et al., 2023). The maximum training round was 500. We used Adam as the optimizer (Kingma & Ba, 2014), with the initial learning rate set to 0.001. If the loss on the validation set did not decrease further within 10 consecutive rounds, the learning rate was halved. When the performance on the validation set did not improve further within 20 consecutive rounds, the training was stopped.
@@ Appendix G Ablation study: Time-frequency Interleaving
Appendix H Visualization

### v2

@@ 3.3.2 Full-frequency-frame attention module
$\boldsymbol{O}_{i}=\text{Softmax}\left(\frac{\boldsymbol{Q}_{i}\boldsymbol{K}_{i}^{\text{T}}}{\sqrt{E\times T}}\right)\boldsymbol{V}_{i}.$ (7)
@@ 3.4 Band restoration module
After going through the separator, the sub-bands need to be converted back to their original width during mask estimation. Specifically, $\boldsymbol{J}_{B}\in\mathbb{R}^{N\times K\times T}$ denotes the output of the separator. For the $k$-th sub-band feature $\boldsymbol{J}_{B,k}\in\mathbb{R}^{N\times T}$ ($k\in[1,K]$), the PReLU activation function and 1D convolutions are used to transform the number of channels to twice the original dimension $2G_{k}$, corresponding to the real and the imaginary part. The complex feature is restored to generate a mask for each sub-band $\boldsymbol{M}_{k}\in\mathbb{C}^{G_{k}\times T}$ using the ReLU activation function. Then they are merged on the frequency dimension to get the mask for the whole band $\boldsymbol{M}\in\mathbb{C}^{F\times T}$. Similar to band-split, the 1D convolutions of different sub-bands do not share parameters.
@@ 4 EchoSet
SoundSpaces 2.0 is an audio rendering platform in 3D environments. Given the mesh of a 3D scenario, it can simulate the acoustic effects of any sound captured from microphones. We followed the steps below to generate mixed speech. (1) Choose the scenario. We selected rooms where daily conversations often occur (such as office, living room, bedroom, dining room, etc.) from Matterport3D, a large RGB-D dataset containing 90 diverse multi-floor and multi-room indoor scenes. (2) Define or sample the position. We defined a microphone at a suitable position, like next to a table or sofa, and sampled two sound sources in the same room. (3) Sample the direction and distance. The angle between the microphone and the sound source must be obtuse, meaning that the speaker and listener face each other. The distance between the microphone and each speaker was randomly sampled between 1 m and 5 m. (4) Sample the height. The microphone and sound sources were randomly generated at a vertical height of 1.5 m to 1.9 m from the floor, which is about a person’s height. (5) Generate the audio. With SoundSpaces 2.0, mixed audio files were generated based on bidirectional path tracking algorithm (Cao et al., 2016), which can simulate various effects in the sound propagation process, including reverberation, diffraction, and absorption. Materials of the room wall and the objects were annotated by Matterport3D and considered during the generation of the audio mixture.
Based on the SoundSpaces 2.0 platform and the Matterport 3D scene dataset, we can simulate reverberant audio from different speakers in LibriSpeech (Panayotov et al., 2015) to build a new dataset, EchoSet. In total, EchoSet includes 20,268 training utterances, 4,604 validation utterances, and 2,650 test utterances. Each utterance lasts for 6 seconds. We mixed the speech of the two speakers at a random overlap ratio and added some noises from WHAM! noise (Wichern et al., 2019). The two different speakers were mixed with signal-to-distortion ratio (SDR) sampled between -5 dB and 5 dB. The noises were mixed with SDR sampled between -10 dB and 10 dB. The dataset is available at: https://huggingface.co/datasets/JusperLee/EchoSet.
@@ 5 Experimental Setup
Dataset. We report the performance of TIGER on EchoSet. For fair comparison with previous speech separation methods (Li et al., 2023; Wang et al., 2023; Hu et al., 2021), we also used two benchmark datasets LRS2-2Mix (Li et al., 2023) and Libri2Mix train-100 min (Cosentino et al., 2020). All of these datasets are at a sampling rate of 16 kHz.
To validate the gap between EchoSet and real-world environments, we constructed real-world data by selecting 10 real-world environments and recording audio from 40 speakers from the LibriSpeech (Panayotov et al., 2015) test set. The two audio used for mixing were recorded in the same acoustic scene (e.g., the shape and material of the walls and objects in the room) and followed the same mixing method as LRS2-2Mix (Li et al., 2023). The duration of each audio is 60 seconds, and the sampling rate is 16 kHz. For more details of these datasets, please refer to Appendix A.
Training and evaluation. During training, we utilized 3-second audio segments for EchoSet and Libri2Mix, and 2-second segments for LRS2-2Mix. The negative SI-SDR was adopted as the training loss (Le Roux et al., 2019). Adam optimizer (Kingma & Ba, 2014) was employed with an initial learning rate of 0.001, adjusted based on validation performance. Evaluation metrics included SDRi and SI-SDRi (Vincent et al., 2006), with higher values indicating better performance. We report parameters and MAC operations for complexity, which are calculated for one second of audio at 16 kHz. Inference speed was measured on NVIDIA RTX 4090 and Intel Xeon Gold 6326. Detailed training and evaluation configurations can be found in Appendix B and Appendix C. Code is available at: https://github.com/JusperLee/TIGER.
@@ 6.1 EchoSet is more close to the real-world data
Figure 4: SI-SDRi results of different models on the real-world data. Models were trained on Libri2Mix, LRS2-2Mix and EchoSet respectively.
@@ 6.2 Comparisons with state-of-the-art methods
Table 2: Performance comparison of TIGER and other separation models on Libri2Mix, LRS2-2Mix, and EchoSet. Models are trained and tested on corresponding datasets. Bold denotes the best performance, and underline indicates the second-best. SDRi and SI-SDRi are recorded in dB.
@@
Separation performance. TIGER obtained competitive separation performance on the three datasets compared with previous SOTA models (see Table 2). On Libri2Mix, which is relatively simple for lack of noise and reverberation, TIGER (large) was second only to the current SOTA model TF-GridNet, with a 6% drop in performance. On LRS2-2Mix, a more complicated dataset with reverberation recorded in different scenes, the drop in performance of TIGER (large) was only 2% compared with TF-GridNet. On EchoSet, the only dataset with the most realistic reverberation among the three, TIGER (large) achieved an SDRi of 14.22 dB, surpassing other existing methods. On this dataset, TIGER (small) also achieved the performance that was only slightly lower than TF-GridNet. From the above experimental results, we can see that the more complex the acoustic scenarios are, the better performance TIGER will produce. Similarly, based on the results in Figure 4, we observed that TIGER also outperforms existing models in real-world test scenarios. This demonstrates that TIGER is applicable to complex real-world acoustic scenarios including diverse noise and reverberation. To visualize the separation result, we present the spectrogram differences between the audio separated by TIGER and TF-GridNet (Appendix I), demonstrating that TIGER is capable of effectively reconstructing both low-frequency and high-frequency features.
TIGER also demonstrates advanced performance on cinematic sound separation, which aims to extract different audio elements from a film’s soundtrack. See Appendix D for details. As for demos of speech and cinematic sound separation, please refer to the project page [^1] https://cslikai.cn/TIGER.
Separation effciency. The parameters of TIGER were only 0.82 M, and the MACs were only 7.65 G/s and 15.27 G/s for the small and large versions respectively. Compared with TF-GridNet, the parameters of TIGER (large) dropped by 94.3%, and the MACs were reduced by 95.3%. For inference of one-second audio, TIGER (large) took about 1/3 of the CPU Time and 3/4 of the GPU Time compared with TF-GridNet, demonstrating a significant calculation compression effect. Besides, TIGER took up less memory during training and inference, making TIGER more suitable for deployment on devices with limited computational resources.
@@ 6.3 Ablation study
We also verified the impact of alternating the frequency path and frame path on model performance, and explored the lightweight potential of TIGER under a smaller parameter scale. See Appendix G and H for more details.
@@ 7 Conclusion
Acknowledgments
This work was supported in part by the National Key Research and Development Program of China (No. 2021ZD0200301) and the National Natural Science Foundation of China (No. U2341228).
@@ Appendix A Dataset details
EchoSet. This dataset includes 20268 training utterances, 4604 validation and 2650 test utterances. The length of each audio is 6 seconds. The target speech was selected from LibriSpeech  (Panayotov et al., 2015), mixed with SDR ranging from -5 dB to 5 dB. The speech and noise which was sampled from WHAM! were mixed with SDR sampled between -10 dB and 10 dB. This dataset contains realistic reverberation. The sampling rate is 16 kHz.
LRS2-2Mix (Li et al., 2023). Each audio in this dataset lasts for 2 seconds, at the sampling rate of 16 kHz. The training set, validation set and test set are about 11.1, 2.8 and 1.7 hours, respectively. The utterances were selected from the LRS2 (Afouras et al., 2018) corpus, which consists of video clips acquired through BBC, and were mixed with SDR sampled between -5 dB and 5 dB. Since the audio files were recorded in real acoustic scenarios, LRS2-2Mix contains much noise and reverberation.
Libri2Mix (Cosentino et al., 2020). Each audio in this dataset lasts for 3 seconds. The target speech for each audio mixture was randomly chosen from LibriSpeech (Panayotov et al., 2015) (train-100) and combined with a uniformly sampled Loudness Units relative to Full Scale (Series, 2011) between -25 and -33 dB. We adoped the 16 kHz min version with no noise or reverberation in our experiments.
Real-world data. We collected a small-scale dataset from the physical world to test the performance of models trained on different datasets in real-world scenarios, with each audio clip 60 seconds long. Its data collection process is described as follows. First, we selected 10 rooms of varying sizes and shapes as distinct acoustic environments. Then, we randomly sampled approximately 1.5 hours of 16 kHz speech audio from the LibriSpeech test set (Panayotov et al., 2015), and sampled noise data from the WHAM! noise dataset (Wichern et al., 2019). During the recording process, audio content was played using the speakers of a 2023 MacBook Pro and recorded via a Logitech Blue Yeti Nano omnidirectional microphone placed in a fixed position. The distance between the speaker and the microphone was randomly selected from 0.3 m to 2 m. The recording parameters were set to a 16 kHz sampling rate and 32-bit depth. This setup ensured that both speech and noise were recorded in the same room, preserving the authenticity of the reverberation effects. Finally, we processed the collected audio by mixing the recordings. Specifically, audio from different speakers was mixed using SDRs randomly sampled between -5 dB and 5 dB. Noise data was added using SDRs randomly sampled between -10 dB and 10 dB. Since the propagation paths of sounds in the air are independent of one another, mixing these components is considered a reasonable approach. This design ensures the realism and diversity of the evaluation dataset, effectively capturing the complexity of speech separation in real-world conditions.
@@ Appendix B Training Configuration
During training, We used a 3-second audio segment for EchoSet and Libri2Mix, and a 2-second for LRS2-2Mix. We used the maximization of SI-SDR as the training loss (Le Roux et al., 2019). The maximum training round was 500. We used Adam as the optimizer (Kingma & Ba, 2014), with the initial learning rate set to 0.001. If the loss on the validation set did not decrease further within 10 consecutive rounds, the learning rate was halved. When the performance on the validation set did not improve further within 20 consecutive rounds, the training was stopped.
@@ Appendix C Evaluation Configuration
When evaluating model performance on real-world data, we used the training lengths (Libri2Mix: 3s, LRS2-2Mix: 2s, EchoSet: 3s) of the respective datasets to inference the 60-second audio with a 50% overlap sliding scale. This approach to some extent mitigates the problem of model performance degradation that may be caused by the difference in training and inference lengths, thus ensuring fairness in the model’s performance comparison on the real-world data.
@@ Appendix G Ablation study: Time-frequency Interleaving
Appendix H The lightweight potential of TIGER
To further illustrate the lightweight potential of TIGER, we present the experimental results of a smaller version of TIGER as well as the compressed SudoRM-RF model (Tzinis et al., 2020) based on GC3 method (Luo et al., 2021) on EchoSet. The hyperparameters of TIGER (tiny) were set as follows: the feature dimension $N$ was reduced from 128 to 24; the hidden layer dimension $H$ was reduced from 256 to 64; the number of FFI blocks $B$ was 4. The hyperparameter settings for SudoRM-RF were strictly in accordance with the publicly available configurations of GC3444https://github.com/yluo42/GC3. The efficiency was evaluated on one-second audio input. As shown in Table 12, TIGER (tiny) significantly outperforms SudoRM-RF + GC3 at comparable efficiencies, which proves that TIGER is a very effective lightweight model.
Method SDRi SI-SDRi Paras MACs GPU Time
(dB) (dB) (K) (G/s) (ms)
SudoRM-RF + GC3 5.0 4.6 303.57 0.81 32.68
TIGER (tiny) 10.7 10.4 102.12 0.72 30.21
Table 12: Performance and efficiency comparison of SudoRM-RF + GC3 and TIGER (tiny).
Appendix I Visualization